How many molecules (not moles) of NH3 are produced from 7.65×10−4 g of H2?
1
Expert's answer
2016-09-20T05:22:46-0400
How many molecules (not moles) of NH3 are produced from 7.65×10−4 g of H2?
Solution:
Production of the NH3 from hydrogen and nitrogen:
3H2 + N2 → 2NH3
3 mol 1 mol 2 mol
The number of moles H2:
n(H2) = m(H2) / M(H2) = 7.65×10−4 g / 2 g/mol = 0.0003825 mol
The number of moles NH3:
n(NH3) = (n(H2) /3) × 2 = (0.0003825 mol / 3 ) × 2 = 0.000255 mol
The number of molecules can be calculated:
N (NH3) = NA × n (NH3) = 6.022 ×1023 mol-1 × 0.000255 mol = 1.53 ×1020 molecules

Answer: 1.53 ×1020 molecules